# Model-Agnostic Bounds for Augmented Inverse Probability Weighted Estimators’ Wald-Confidence Interval Coverage in Randomized Controlled Trials

### v1

@@ 6 Discussion
Taking $\delta=({\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2})^{1/3}$ yields the desired rate. Additionally under subgaussian conditions 10 and S2, $\tilde{\Delta}(\delta)$ can be replaced by $\tilde{\Delta}^{\prime}(\delta)$ if $2\delta M>{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}$. Taking $\delta=\{{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}+\tilde{s}_{1}\sqrt{2{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}\log n}\}/(2M)$, which tends to zero as $n\to\infty$, yields the other desired rate.
@@
$J(2\delta,{\mathcal{F}},M)\leq\int_{0}^{2\delta}\sqrt{C_{2}(\nu)\epsilon^{-2+2/\nu}}{\mathrm{d}}\epsilon=C_{2}(\nu)\int_{0}^{2\delta}\epsilon^{-1+1/\nu}{\mathrm{d}}\epsilon=C_{2}(\nu)\delta^{1/\nu}.$ (S14)
@@
${\mathrm{O}}\left(\tilde{\Delta}(\delta)+\delta^{1/\nu}+\frac{1}{\delta^{2-2/\nu}\sqrt{n}}+(\delta+n^{-1/2})\log n\right).$
Taking $\delta=({\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2})^{\nu/(1+2\nu)}+n^{-\nu/(4\nu-2)}$ yields the desired rate. Additionally under Conditions 10 and S2, taking $\delta=\{{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}+\tilde{s}_{1}\sqrt{2{\mathbb{E}}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2}\log n}\}/(2M)$, which tends to zero as $n\to\infty$, yields the other desired rate. ∎
@@
$={\mathbb{E}}[P_{n}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-\tilde{\psi}_{a}^{2}]-P_{*}{\mathcal{T}}_{a}(Q_{\#,a})^{2}+\psi_{*,a}^{2}$
$={\mathbb{E}}\left[P_{n}\left\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\right\}\right]-{\mathrm{Var}}(\tilde{\psi}_{a})$
$={\mathbb{E}}\left[P_{n}\left\{\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)^{2}}\right\}\left\{[Y-\tilde{Q}_{a}(X)]^{2}-[Y-Q_{\#,a}(X)]^{2}\right\}\right]$
$\quad+2{\mathbb{E}}\left[P_{n}\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)}\{Y-\tilde{Q}_{a}(X)\}\tilde{Q}_{a}(X)\right]-2P_{*}\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)}\{Y-Q_{\#,a}(X)\}Q_{\#,a}(X)$
$\quad+{\mathbb{E}}[P_{n}(\tilde{Q}_{a}^{2}-Q_{\#,a}^{2})]-{\mathrm{Var}}(\tilde{\psi}_{a}).$

### v2

@@ 6 Discussion
Taking $\delta=({E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2})^{1/3}$ yields the desired rate. Additionally under approximate sub-Weibull conditions 11 and 12, $\tilde{\Delta}(\delta)$ can be replaced by $\tilde{\Delta}^{\prime}(\delta)$. Taking $\delta=\tilde{r}_{1}(n)[(1+\tilde{d}_{1})\tilde{c}_{1}^{-1}\log n]^{1/\tilde{q}_{1}}/(2M)$, which tends to zero as $n\to\infty$, yields the other desired rate.
@@
$J(2\delta,{\mathcal{F}},M)\leq\int_{0}^{2\delta}\sqrt{C_{2}(\nu)\epsilon^{-2\nu/(2\nu+1)}}{\mathrm{d}}\epsilon=C_{2}(\nu)\int_{0}^{2\delta}\epsilon^{-\nu/(2\nu+1)}{\mathrm{d}}\epsilon=C_{2}(\nu)\delta^{(\nu+1)/(2\nu+1)}.$ (S14)
@@
${\mathrm{O}}\left(\tilde{\Delta}(\delta)+\delta^{(\nu+1)/(2\nu+1)}+\frac{1}{\delta^{2\nu/(2\nu+1)}\sqrt{n}}+(\delta+n^{-1/2})\log n\right).$
Taking $\delta=({E}\|\tilde{Q}_{a}-Q_{\#,a}\|_{2}^{2})^{(2\nu+1)/(5\nu+3)}+n^{-(2\nu+1)/(6\nu+2)}$ yields the desired rate. Additionally under Conditions 11 and 12, taking $\delta=\tilde{r}_{1}(n)[(1+\tilde{d}_{1})\tilde{c}_{1}^{-1}\log n]^{1/\tilde{q}_{1}}/(2M)$, which tends to zero as $n\to\infty$, yields the other desired rate. ∎
@@
$={E}[P_{n}{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-\tilde{\psi}_{a}^{2}]-P_{*}{\mathcal{T}}_{a}(Q_{\#,a})^{2}+\psi_{*,a}^{2}$
$={E}\left[P_{n}\left\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\right\}\right]-{\mathrm{var}}(\tilde{\psi}_{a})$
$={E}\left[P_{n}\left\{\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)^{2}}\right\}\left\{[Y-\tilde{Q}_{a}(X)]^{2}-[Y-Q_{\#,a}(X)]^{2}\right\}\right]$
$\quad+2{E}\left[P_{n}\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)}\{Y-\tilde{Q}_{a}(X)\}\tilde{Q}_{a}(X)\right]-2P_{*}\frac{{\mathbbm{1}}(A=a)}{\pi_{*}(a\mid X)}\{Y-Q_{\#,a}(X)\}Q_{\#,a}(X)$
$\quad+{E}[P_{n}(\tilde{Q}_{a}^{2}-Q_{\#,a}^{2})]-{\mathrm{var}}(\tilde{\psi}_{a}).$
@@
S5.4 Series regression
I will use the notation in Section S2.
Proof of Proposition S1.
Since regressions with $T_{K}\Phi_{K}(X)$ and $\Phi_{K}(X)$ lead to the same estimators, without loss of generality, assume that $T_{K}$ is identity. By SeriesCond S2, $P_{*}\Phi_{K}(X)\Phi_{K}(X)^{\top}$ is invertible. Let $Z:=\Phi_{K}(X)\Phi_{K}(X)^{\top}-P_{*}\Phi_{K}(X)\Phi_{K}(X)^{\top}$ and $\mathcal{E}$ denote the event that $\lambda_{\max}((P_{n}-P_{*})\Phi_{K}(X)\Phi_{K}(X)^{\top})\leq\underline{\lambda}/2$. By SeriesCond S1 and S2, $\lambda_{\max}(Z)\leq\lambda_{\max}(\Phi_{K}(X)\Phi_{K}(X)^{\top})+\lambda_{\max}(P_{*}\Phi_{K}(X)\allowbreak\Phi_{K}(X)^{\top})\leq\xi_{K}^{2}+\bar{\lambda}$. Since ${E}[Z^{2}]=P_{*}\Phi_{K}(X)\Phi_{K}(X)^{\top}\Phi_{K}(X)\Phi_{K}(X)^{\top}-(P_{*}\Phi_{K}(X)\Phi_{K}(X)^{\top})^{2}$ and $\lambda_{\max}(P_{*}\Phi_{K}(X)\Phi_{K}(X)^{\top}\Phi_{K}(X)\Phi_{K}(X)^{\top})\leq\xi_{K}^{2}\lambda_{\max}(P_{*}\Phi_{K}(X)\Phi_{K}(X)^{\top})\leq\xi_{K}^{2}\bar{\lambda}$, it holds that $\lambda_{\max}(Z^{2})\leq\xi_{K}^{2}\bar{\lambda}$. Applying matrix Bernstein inequality (e.g., Theorem 6.1.1 in Tropp,, 2015) to the sample mean of $Z$, namely $P_{n}Z=(P_{n}-P_{*})\Phi_{K}(X)\Phi_{K}(X)^{\top}$, yields that
${\mathrm{pr}}(\mathcal{E})\geq 1-2K\exp\left\{-\frac{n\underline{\lambda}^{2}}{8\xi_{K}^{2}\bar{\lambda}+4(\xi_{K}^{2}+\bar{\lambda})\underline{\lambda}/3}\right\}.$
By Weyl’s inequality, when $\mathcal{E}$ occurs, $\lambda_{\min}(P_{n}\Phi_{K}(X)\Phi_{K}(X)^{\top})\geq\underline{\lambda}-\lambda_{\max}((P_{n}-P_{*})\Phi_{K}(X)\Phi_{K}(X)^{\top})\geq\underline{\lambda}/2>0$ and so $P_{n}\Phi_{K}(X)\Phi_{K}(X)^{\top}$ is invertible. Moreover, $\lambda_{\max}(P_{n}\Phi_{K}(X)\Phi_{K}(X)^{\top})\leq\lambda_{\max}(P_{*}\Phi_{K}(X)\allowbreak\Phi_{K}(X)^{\top})+\lambda_{\max}((P_{n}-P_{*})\Phi_{K}(X)\Phi_{K}(X)^{\top})\leq\bar{\lambda}+\underline{\lambda}/2$.
Let $\epsilon:=Y-Q_{\#}(X)=Y-\Phi_{K}(X)^{\top}\beta_{\#}$ and $W:=\Phi_{K}(X)\epsilon$. Since $Y$ is bounded and $Q_{\#}$ is bounded by some constant independent of $K$ when $\mathcal{E}$ occurs. Thus, when $\mathcal{E}$ occurs, $\epsilon$ is uniformly bounded in $K$. Suppose that $\mathcal{E}$ occurs, then
$\hat{\beta}-\beta_{\#}=\{P_{n}\Phi_{K}(X)\Phi_{K}(X)^{\top}\}^{-1}\{P_{n}\Phi_{K}(X)Y-P_{n}\Phi_{K}(X)\Phi_{K}(X)^{\top}\beta_{\#}\}=\{P_{n}\Phi_{K}(X)\Phi_{K}(X)^{\top}\}^{-1}P_{n}W.$
and
$\|\hat{Q}-Q_{\#}\|_{2}^{2}$ $=P_{*}\{\Phi_{K}(X)^{\top}(\hat{\beta}-\beta_{\#})\}^{2}$
$=(\hat{\beta}-\beta_{\#})^{\top}\{P_{*}\Phi_{K}(X)\Phi_{K}(X)^{\top}\}(\hat{\beta}-\beta_{\#})$
$=\{P_{n}W^{\top}\}\{P_{n}\Phi_{K}(X)\Phi_{K}(X)^{\top}\}^{-1}\{P_{*}\Phi_{K}(X)\Phi_{K}(X)^{\top}\}\{P_{n}\Phi_{K}(X)\Phi_{K}(X)^{\top}\}^{-1}\{P_{n}W\}.$
By construction, $P_{*}W=0$. Applying matrix Bernstein inequality to $P_{n}W$ yields that, there exists a constant $C_{2}$ independent of $(K,n)$ such that, for any $t>0$,
${\mathrm{pr}}(\|P_{n}W\|_{\ell_{2}}>t)\leq(K+1)\exp\left\{-C_{2}\frac{nt^{2}}{2\xi_{K}^{2}+2\xi_{K}t/3}\right\}.$
Thus, for any $t>0$,
${\mathrm{pr}}\left(\frac{\|\hat{Q}-Q_{\#}\|_{2}}{\sqrt{K/n}}>t\right)={\mathrm{pr}}\left(\frac{\|\hat{Q}-Q_{\#}\|_{2}^{2}}{K/n}>t^{2}\right)$
$\leq{\mathrm{pr}}\left(\frac{\|\hat{Q}-Q_{\#}\|_{2}^{2}}{K/n}>t^{2},\mathcal{E}\right)+1-{\mathrm{pr}}(\mathcal{E})$
$\leq{\mathrm{pr}}\left(\frac{n}{K}\frac{\bar{\lambda}}{(\underline{\lambda}/2)^{2}}\|P_{n}W\|_{{\ell_{2}}}^{2}>t^{2}\right)+2K\exp\left\{-\frac{n\underline{\lambda}^{2}}{8\xi_{K}^{2}\bar{\lambda}+4(\xi_{K}^{2}+\bar{\lambda})\underline{\lambda}/3}\right\}$
$\leq(K+1)\exp\left\{-A_{1}\frac{\sqrt{n}Kt^{2}}{A_{2}\sqrt{n}\xi_{K}^{2}+A_{3}\xi_{K}\sqrt{K}t}\right\}+2K\exp\left\{-\frac{n\underline{\lambda}^{2}}{8\xi_{K}^{2}\bar{\lambda}+4(\xi_{K}^{2}+\bar{\lambda})\underline{\lambda}/3}\right\}$
where $A_{1},A_{2},A_{3}$ are constants independent of $(K,n)$. The desired result follows from SeriesCond S3. ∎
It is possible to use concentration inequalities for vector $\ell_{2}$ norms (e.g. Pinelis,, 1994; Maurer and Pontil,, 2021), particularly those centered at the expected norm, to obtain sharper tail bounds.